Hamilton Company uses a periodic inventory system. At the end of the annual accounting period, December 31 of the current year, the accounting records provided the following information for product 1: Units Unit Cost Inventory, December 31, prior year 1,930 $ 6 For the current year: Purchase, March 21 6,010 5 Purchase, August 1 4,120 3 Inventory, December 31, current year 2,900 Required: Compute ending inventory and cost of goods sold under FIFO, LIFO, and average cost inventory costing metho
Answer #1
Unit Unit cost Total
Beginning inventory 1930 6 11580
Purchase march 21 6010 5 30050
Purchase august 1 4120 3 12360
Total 12060 53990

Calculate ending inventory and cost of goods sold

FIFO LIFO Average cost
Cost of ending inventory 2900*3 = 8700 (1930*6+970*5) = 16430 53990/12060*2900 = 12983
Cost of goods sold 53990-8700 = 45290 53990-16430 = 37560 53990-12983 = 41007